Use Gauss’ Law to calculate the electric field due to an infinite non-conducting sheet of charge, with surface charge density 


1
Expert's answer
2021-04-29T10:44:35-0400

Answer

Consider a non conducting sheet with charge density "\sigma"


By gauss law flux of it

"\phi=\frac{\sigma s}{\epsilon_0}"


And by definition

"\phi=\int Edscos0°+\int Eds cos0°"

So


"\frac{\sigma s}{\epsilon_0}=2Es\\E=\frac{\sigma }{2\epsilon_0}"